what is
5x+1=x+13

Answer by funmath(2933)

This is an equation.
:
5x+1=x+13
5x-x+1=x-x+13
4x+1=13
4x+1-1=13-1
4x=12
4x/4=12/4
x=3
:
Check, if x=3, does the equation balance?
5(3)+1=3+13
15+1=16
16=16 We're right!
Happy Calculating!!!